OpenStudy (anonymous):

\[3\sqrt{2}+\sqrt{8}\]

OpenStudy (anonymous):

@alex1248191

OpenStudy (anonymous):

answer is 5 sqr of 2!

OpenStudy (kirbykirby):

\(\sqrt{8}=\sqrt{4\cdot 2}=\sqrt{4}\sqrt{2}=2\sqrt{2}\)

OpenStudy (kirbykirby):

then you can add the two terms together by just treating \(\sqrt{2}\) as if it were a variable

OpenStudy (anonymous):

thank you @kirbykirby and @alex1248191 :)

OpenStudy (kirbykirby):

=]